OpenStudy (anonymous):

REALLY NEED HELP QUICK! Simplify (4x/x^2-9)/(8x^2/x^2+6x+9)

OpenStudy (callisto):

\[\frac{\frac{4x}{x^2-9}}{\frac{8x^2}{x^2+6x+9}}=\frac{4x}{x^2-9}\times \frac{x^2+6x+9}{8x^2}=\frac{4x}{(x-3)(x+3)}\times \frac{(x+3)^2}{8x^2}\]\[=\frac{1}{(x-3)}\times \frac{(x+3)}{2x} =?\]

OpenStudy (anonymous):

x+3/2x(x-3)?

OpenStudy (stacey):

Yes. That is correct.